# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

### v2

@@ 4.2 The Orientation of the Magnetic Field Throughout the CMZ
Building on the comparison of magnetic field scales presented in Section 4.1, we study the distribution of projected polarization pseudovectors to determine systematic trends in the polarization angle (and by extension the magnetic field) throughout the CMZ. The objective here is to quantify the extent to which the magnetic field derived from FIREPLACE connects to the large-scale magnetic field observed by PILOT and ACTPol. We plot the total intensity versus polarization angle in Figure 12, where the individual data points are colored based on the polarization percentage (/100) for each line-of-sight. The polarization angle distribution covers the full range of angles from -90 to +90. Furthermore, an inspection of the observed polarization percentage of the observations reveals that polarization levels 20% are rare and only present at 3,000 MJy sr. Only measurements have a polarization percentage 20%, comprising 1% of our entire data set. These high polarization measurements are likely a result of systematic effects or statistical fluctuations. The 3000 MJy sr regime corresponds to the peripheries of the molecular clouds in the CMZ, revealing that the polarization percentage is generally higher in regions of lower total emission. This emission regime is demarcated at its upper bound by the lower black horizontal dashed line shown in Figure 12.
At the highest total intensities (30,000 MJy sr) the number of polarization angles obtained is more sparse. This total intensity regime corresponds to the brightest molecular clouds in the CMZ, including Sgr B2, Sgr C, and the Brick. The fractional polarizations are generally 5% in this regime. This high total intensity regime is demarcated at its lower bound by the upper horizontal dashed line in Figure 12. The generally lower polarization percentages in these denser regions relative to the percentages in the fainter cloud peripheries agrees with previous polarimetric results for CMZ molecular clouds (e.g. Chuss et al. 2003a, b; Pillai et al. 2015; Lu et al. 2023; FIREPLACE I).
The large number of individual lines of sight in Figure 12 (65,000 Nyquist-sampled pseudovectors) makes it difficult to assess whether there are any preferential polarization angle orientations in the CMZ. To obtain better insight into CMZ-wide polarization angle trends we sort the individual polarization angles shown in Figure 12 into 10 bins. The histogram representation of the binned polarization angles is shown in Figure 13 (similar to what is done in FIREPLACE I). In these histograms, an angle of 0 indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90 indicating a magnetic field that is oriented perpendicular to the Galactic plane.
@@
Column A illustrates the fractional polarization distribution for vectors of each intensity range. We observe a steeper slope for the higher total intensities (where 30,000 MJy sr). This steeper distribution was also observed in FIREPLACE I and could indicate increased depolarization towards the brighter (higher density) regions. This increased depolarization could be a result of decreased grain alignment efficiency in the interiors of molecular clouds (Santos et al., 2019), superpositions of multiple field components along the line-of-sight, or field tangling on scales smaller than the 19.6″ SOFIA/HAWC+ beam size of these observations (e.g. Fissel et al., 2016). Higher resolution observations are needed to determine which of these mechanisms is responsible for the increased depolarization observed in our results. We see an increasingly shallow slope as we move to fainter intensity regimes as marked in the third and fourth rows of Figure 13. This trend agrees with what was observed in DR1 (FIREPLACE I).
@@
The histograms presented in columns B - D reveal preferential polarization angle orientations. The histogram corresponding to all of the polarization pseudovectors (top row) reveals a bimodal distribution in the polarization angle, where the histogram peaks at an angle of approximately -25 and 90. This is a similar bimodal distribution to the one observed in FIREPLACE I. We observe the same polarization angle enhancements in the middle and low emission regimes as seen in rows 3 and 4 of Figure 13. In the high regime presented in the second row of Figure 13 there is no 90 enhancement and there is instead an enhancement at 0.
@@
The peak at in the brightest intensity regime that corresponds to the emission originating from the central, brightest and densest regions of the prominent CMZ molecular clouds. Since this distribution is connected to high emission () regions we find it unlikely that the orientation at is a magnetic field component external to the CMZ.
@@
We also bin the 2022 polarization angle distribution shown in Figure 3 to assess whether the same polarization angle enhancements are recovered in this portion of the FIREPLACE observations. These histograms are presented in Appendix A.2 and shown in Figure A.1. We find a similar bimodal polarization angle enhancement at angles of -20 and -80 in all 2022 pseudovectors and in the middle and low emission regimes. These are similar bimodal enhancements seen in the full FIREPLACE data set in these intensity regimes. However, in the high emission regime we do not detect the 0 enhancement and instead observe an enhancement at -60. Therefore, the 0 component seen in the emission regime of the full data set can be associated with Sgr B2. Since this high intensity regime corresponds to the cores of the brightest and most dense molecular clouds, the different enhancements observed in this regime indicate that the magnetic field is preferentially aligned along the morphology of these clouds, further corroborating that the magnetic field in this regime is local to the CMZ.
@@
To further quantify the bimodal distribution from the full FIREPLACE data set, we fit the histogram of polarization pseudovectors obtained for the full FIREPLACE data set with a two-Gaussian model. To perform this fitting, we assume the uncertainties on the counts in each histogram bin is the square-root of the counts in that bin (ranging from 44100 counts). These fits are shown in Figure 14, where the polarization angles derived from FIREPLACE have been sorted into 5 bins for the full set of polarization angles and the middle intensity regime. The blue line is the best-fit result obtained from fitting a two-Gaussian model to the histogram of the polarization angle data defined as:
@@
where is the offset above zero for the Gaussian components (in counts), and are the amplitudes of each Gaussian (in counts), is the set of polarization angles from -90 to +90 that comprise the histogram, and are the centroids of each Gaussian (in ), and and are the standard deviations of each Gaussian (in ). This results in a seven-parameter model, with all eight parameters allowed to vary to determine the best-fit to the underlying histogram. The resulting best fit parameters for the full polarization vector distribution and the middle intensity regime are shown in Table 1.
@@
The peaks of the modelled Gaussian components are located at polarization angles of and . The residuals are shown below the model fit, and range from a few hundred to 1000 counts. These residuals are fairly large compared to the uncertainties in the histogram bins, but this two-Gaussian model fit is simple, and is likely not fully characterizing the structure of the histogram.
@@
The histogram and resulting best-fit model for the middle intensity regime corresponding to the third row of Figure 13 is shown in the right panel of Figure 14. The fit returns similar parameters that are consistent within the error bars of the two measurements with peaks located at and . As with the full polarization angle pseudovector, the residuals corresponding to the best-fit model are implying that the model is under-fitting the structure of the histogram. More detailed characterization of these histograms would require a model consisting of additional model components. Theoretical modeling is needed to estimate possible additional model components.
@@
The polarization angle peak from the model fit at approximately corresponds to a magnetic field orientation that is aligned with the CMZ-wide magnetic field observed by the ACTPol and PILOT programs (Guan et al., 2021; Mangilli et al., 2019). As discussed this field is connected to material in the inner 12 kpc of the Galaxy (Nishiyama et al., 2009). This enhancement could therefore be tracing the orientation of the CMZ molecular clouds or it could be a magnetic field component external to the CMZ. The Gaussian component centered at approximately relates, conversely, to a vertical magnetic field and provides evidence of a connection between the magnetic fields in the dust to those traced by the radio NTF population. We do not fit a two-Gaussian model to the high-intensity regime because of the lower number of polarization angles per histogram bin in this regime as can be seen in Figure 13 – 100 compared to the full distribution with 1000.
@@
The high-total intensity regime cutoff of 310 MJy sr is lower than was used in FIREPLACE I, where a threshold of 510 MJy sr was used to isolate the magnetic field orientation obtained from only the brightest regions of Sgr B2. As a further verification of the combined FIREPLACE data set we analyze the histogram of polarization angles obtained from using the DR1 upper total intensity regime of 510 MJy sr to isolate the polarization angles from Sgr B2. We would expect to obtain the same distribution as shown in FIREPLACE I. We find the same number of polarization pseudovectors to what was found in DR1 (359) corresponding to this regime where 510. We find essentially the same histogram distribution as reported in Figure 8 of FIREPLACE I. There is some variation where certain 10 bins are one count higher or lower in our combined observations compared to the histogram presented in FIREPLACE I. Such slight differences in individual histogram bins are likely a result of slight rounding differences where polarization angles near the 10 bin edges are placed in different bins.
@@
The orientation of the magnetic field enhancements help to clarify a picture of the magnetic fields across the central degree of the Galaxy. In all histograms, we see excess measurements at an angle consistent with that of the large-scale component seen in the PILOT and ACTPol data (the -20 enhancement discussed for Figures 13, 14, A.1). In the high intensity regime (the yellow histogram of Figure 13), the polarization angle enhancements observed at -20 splits into two sub-components: one at -20 and one at 0. However, in the 2022 high intensity portion shown in Figure A.1 the enhancements are at -60 and -20.
@@
The 0enhancement is only observed in the brightest portion of Sgr B2 is almost certainly associated with the core of the Sgr B2 region, since we do not recover this component in the 2022 observations (Figure A.1). The -20 component is detected in all our intensity regimes. In the high intensity regime the -20 enhancement is likely tracing a local CMZ field. For the lower intensity regimes, it could be at least partially an external magnetic field component given the large spatial extent over which this field is coherent at arcminute resolutions. The enhancement at 90 is only detected in the middle and low intensity regimes. The detection of this enhancement in only these regimes is consistent with the picture of a connection between the vertical field traced by the NTFs and a horizontal field that is a result of the field being sheared in denser, kinematically-dominated regions.
@@
Previously, it has been posited that an initially poloidal field could be sheared into a horizontal configuration in denser molecular regions (Morris & Serabyn, 1996). Our findings here support early submillimeter data (Novak et al., 2003; Chuss et al., 2003a) and near-infrared polarimetry data (Nishiyama et al., 2010) that have suggested a transition occurs from a sheared field in dense regions to a vertical field in regions having lower kinetic energy density. The two scenarios that have been considered are the above-mentioned evolution from an initially poloidal field and the creation of a vertical field from the sheared field via winds and outflows.
@@
In producing these histograms, we have utilized polarimetry data across the range of intensities down to 1000 MJy sr. If systematic errors remain after our data cuts, they are unlikely to produce the histogram peaks observed. The scan angles () are sufficiently different from the maxima of the histograms that a systematic scan-synchronous effect is unlikely to produce the results seen here.
| Parameter | All Polarization Angles | Middle Intensity Regime |
| (counts) | 3810 | 1850 |
| (counts) | 5780 | 4250 |
| (counts) | 6720 | 5450 |
| () | -82.1 | -84.3 |
| () | -18.6 | -19.6 |
| () | 17.7 | 16.4 |
| () | 22.4 | 24.4 |

### v3

@@ 4.2 The Orientation of the Magnetic Field Throughout the CMZ
Building on the comparison of magnetic field spatial scales presented in Section 4.1, we study the distribution of polarization pseudovectors to determine systematic trends in the polarization angle (and by extension the magnetic field) throughout the CMZ. The objective here is to quantify the extent to which the magnetic field derived from FIREPLACE connects to the large-scale magnetic field observed by PILOT and ACTPol. We plot the total intensity versus polarization angle in Figure 12, where the individual data points are colored based on the polarization percentage (/100) for each line-of-sight. The polarization angle distribution covers the full range of angles from -90∘ to +90∘. Furthermore, an inspection of the observed polarization percentage of the observations reveals that polarization levels 20% are rare (1% of our total pseudovectors).
The horizontal dashed lines divide the plot into different total intensity regimes. The 3000 MJy sr-1 regime corresponds to the emission level of the peripheries of the molecular clouds in the CMZ, whereas 30,000 MJy sr-1 corresponds to the brightest molecular emission in the CMZ of the most prominent CMZ clouds (e.g., Sgr B2, Sgr C, and the Brick molecular clouds).
The higher polarization percentages of 20% are only present in the fainter emission regime, corresponding to only 100 lines-of-sight. These high polarization measurements connect to the previously observed trend that the polarization percentage is generally higher in the fainter peripheries of CMZ molecular clouds. These high values could also indicate large-scale, low-surface brightness emission is missing in our observations due to the OTFMAP mode.
At the highest total intensities (30,000 MJy sr-1, demarcated by the upper horizontal dashed line in Figure 12) the number of polarization angles obtained is more sparse. The polarizations are generally 5% in this regime. The generally lower polarization percentages in these denser regions relative to the percentages in the fainter cloud peripheries agrees with previous polarimetric results for CMZ molecular clouds (e.g. Chuss et al. 2003a, b; Pillai et al. 2015; Lu et al. 2024; FIREPLACE I).
The large number of individual lines-of-sight in Figure 12 (64,000 Nyquist-sampled pseudovectors) makes it difficult to assess whether there are any preferential polarization angle orientations in the CMZ. To obtain better insight into CMZ-wide polarization angle trends we sort the individual polarization angles shown in Figure 12 into 10∘ bins. The histogram representation of the binned polarization angles is shown in Figure 13 (similar to what is done in FIREPLACE I). In these histograms, a polarization angle of 0∘ indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90∘ indicating a magnetic field that is oriented perpendicular to the Galactic plane.
@@
Column A illustrates the fractional polarization distribution for pseudovectors of each intensity range. We observe a steeper slope for the higher total intensities (where 30,000 MJy sr-1). This steeper distribution was also observed in FIREPLACE I and could indicate increased depolarization towards the brighter (higher density) regions. This increased depolarization could be a result of decreased grain alignment efficiency in the interiors of molecular clouds (Santos et al., 2019), superpositions of multiple field components along the line-of-sight, or field tangling on scales smaller than the 19.6″ SOFIA/HAWC+ beam size of these observations (e.g. Fissel et al., 2016). Higher resolution observations are needed to determine which of these mechanisms is responsible for the increased depolarization observed in our results. We see an increasingly shallow slope as we move to fainter intensity regimes as marked in the third and fourth rows of Figure 13. This trend agrees with what was observed in DR1 (FIREPLACE I).
@@
The histograms presented in columns B - D reveal preferential polarization angle orientations. The histogram corresponding to all of the polarization pseudovectors (top row) reveals a bimodal distribution in the polarization angle, where the histogram peaks at an angle of approximately -20∘ and 90∘. This is a similar bimodal distribution to the one observed in FIREPLACE I. We observe the same polarization angle enhancements in the middle and low emission regimes as seen in rows 3 and 4 of Figure 13. In the high regime presented in the second row of Figure 13 there is no 90∘ enhancement and there is instead an enhancement at 0∘.
@@
The peak at ∘ only appears in the brightest intensity regime corresponding to the emission originating from the brightest and densest regions of the prominent CMZ molecular clouds. We therefore find it is unlikely that the orientation at ∘ is a magnetic field component external to the CMZ.
@@
We also bin the 2022 polarization angle distribution shown in Figure 3 to assess whether the same polarization angle enhancements are recovered in the DR2 portion of the FIREPLACE observations. These histograms are presented in Appendix A.3 and shown in Figure A.2. We find a similar bimodal polarization angle enhancement at angles of -20∘ and -70∘ in all 2022 pseudovectors and in the middle and low emission regimes. These are similar bimodal enhancements to those observed in the full FIREPLACE data set shown in Figure 13 for these intensity regimes. However, in the high emission regime of only the 2022 portion of our FIREPLACE data set we do not detect the 0∘ enhancement and instead observe an enhancement at -60∘, as seen in Figure A.2. Therefore, the 0∘ component seen in the emission regime of the full data set can be associated with Sgr B2. Since this high intensity regime corresponds to the cores of the brightest and most dense molecular clouds, the different enhancements observed in this regime indicate that the magnetic field is preferentially aligned along the morphology of these clouds, further corroborating the conclusion that the magnetic field in this regime is local to the CMZ molecular clouds.
@@
A more comprehensive analysis of the magnetic field alignment will be conducted in a further work utilizing the Histogram of Relative Orientation (HRO, Soler et al., 2013) and Projected Rayleigh Statistic (PRS, Jow et al., 2018) methods. For now, to further quantify the bimodal distribution from the full FIREPLACE data set, we fit the histogram of polarization pseudovectors obtained for the full FIREPLACE data set with a two-Gaussian model. To perform this fitting, we assume the uncertainties on the counts in each histogram bin is the square-root of the counts in that bin (ranging from 3151 counts). These fits are shown in Figure 14, where the polarization angles derived from FIREPLACE have been sorted into 5∘ bins for the full set of polarization angles and the middle intensity regime. The histogram and model have been shifted by 60∘ in the figure to better show the bimodal distribution in the histogram and resulting fit. We note that though the range of angles is technically wrapped such that 60∘ connects to 59∘, the model fit was performed on a linear representation of our polarization angles where 60∘ and 59∘ are on opposite sides of the histogram as shown in Figure 14. This disconnect in linear space is why the model fit is not continuous when wrapped. The blue line is the best-fit result obtained from fitting a two-Gaussian model to the histogram of the polarization angle data with the shadowed envelope showing the 1 dispersion of the model fit. The equation of this two-Gaussian model is defined as:
@@
where is the offset above zero for the Gaussian components (in counts), and are the amplitudes of each Gaussian (in counts), is the set of polarization angles from -90∘ to +90∘ that comprise the histogram, and are the centroids of each Gaussian (in ∘), and and are the standard deviations of each Gaussian (in ∘). This results in a seven-parameter model, with all parameters allowed to vary to determine the best-fit to the underlying histogram. The resulting best fit parameters for the full polarization vector distribution and the middle intensity regime are shown in Table 1 with 1 uncertainties.
@@
The peaks of the modelled Gaussian components for the full set of pseudovectors are located at polarization angles of∘ and ∘. The residuals are shown below the model fit, and range from 10 to a few hundred counts. These residuals are fairly large compared to the uncertainties in the histogram bins, but this two-Gaussian model fit is simple, and is likely not fully characterizing the structure of the histogram.
@@
The histogram and resulting best-fit model for the middle intensity regime corresponding to the third row of Figure 13 is shown in the right panel of Figure 14. The fit returns similar parameters that are generally consistent within the error bars of the Gaussian model components found for the full pseudovector distribution. The Gaussian peaks for the middle intensity regime are located at ∘ and ∘. As with the full polarization angle model fit, the residuals corresponding to the best-fit model are implying that the model is under-fitting the structure of the histogram. More detailed characterization of these histograms would require a model consisting of additional model components. Theoretical modeling is needed to estimate possible additional model components.
@@
The polarization angle peak from the model fit at approximately ∘ corresponds to a magnetic field orientation that is aligned with the CMZ-wide magnetic field observed by the PILOT and ACTPol programs (Mangilli et al., 2019; Guan et al., 2021). As discussed this field is connected to material in the inner 12 kpc of the Galaxy (Nishiyama et al., 2009). This enhancement could therefore be tracing the orientation of the CMZ molecular clouds or it could be a magnetic field component external to the CMZ. The Gaussian component centered at approximately ∘ relates, conversely, to a vertical magnetic field and provides evidence of a connection between the magnetic fields in the dust to those traced by the radio NTF population. We do not fit a two-Gaussian model to the high-intensity regime because of the lower number of polarization angles per histogram bin in this regime as can be seen in Figure 13 – 100 counts compared to the full distribution with 1000 counts.
@@
The high-total intensity regime cutoff of 3104 MJy sr-1 is lower than was used in FIREPLACE I, where a threshold of 5104 MJy sr-1 was used to isolate the magnetic field orientation obtained from only the brightest regions of Sgr B2. As a further verification of the combined FIREPLACE data set we analyze the histogram of polarization angles obtained from using the DR1 upper total intensity regime of 5104 MJy sr-1 to isolate the polarization angles from Sgr B2. We would expect to obtain the same distribution as shown in FIREPLACE I. We find the same number of polarization pseudovectors to what was found in DR1 (359) corresponding to this regime where 5104. We find essentially the same histogram distribution as reported in Figure 8 of FIREPLACE I. There is some variation where certain 10∘ bins are one count higher or lower in our combined observations compared to the histogram presented in FIREPLACE I. Such slight differences in individual histogram bins are likely a result of slight rounding differences where polarization angles near the 10∘ bin edges are placed in different bins.
@@
The orientation of the magnetic field enhancements helps to clarify a picture of the magnetic fields across the central degree of the Galaxy. In all histograms, we see excess measurements at an angle consistent with that of the large-scale component seen in the PILOT and ACTPol data (the -20∘ enhancement shown in Figures 13, 14, A.2).
@@
In the high intensity regime (the yellow histogram of Figure 13), the polarization angle enhancements observed at -20∘ splits into two sub-components: one at -20∘ and one at 0∘. However, in the 2022 high intensity portion shown in Figure A.2 the enhancements are at -60∘ and -20∘. The 0∘ enhancement is only observed in the brightest portion of Sgr B2 is almost certainly associated with the core of the Sgr B2 region, since we do not recover this component in the 2022 observations (Figure A.2). The -20∘ component is detected in all our intensity regimes. In the high intensity regime the -20∘ enhancement is likely tracing a local CMZ field. For the lower intensity regimes, it could be at least partially an external magnetic field component given the large spatial extent over which this field is coherent at arcminute resolutions (e.g. ACTPol, Guan et al., 2021). The enhancement at 90∘ is only detected in the middle and low intensity regimes. The detection of this enhancement in only these regimes is consistent with the picture of a connection between the vertical field traced by the NTFs and a horizontal field that is a result of the field being sheared in denser, kinematically-dominated regions.
@@
Previously, it has been posited that an initially vertical (or poloidal) field could be sheared into a horizontal configuration in denser molecular regions (Morris & Serabyn, 1996). Our findings here support early submillimeter data (Novak et al., 2003; Chuss et al., 2003a) and near-infrared polarimetry data (Nishiyama et al., 2010) that have suggested a transition occurs from a sheared field in dense regions to a vertical field in regions having lower kinetic energy density. The two scenarios that have been considered are the above-mentioned evolution from an initially vertical field and the creation of a vertical field from the sheared field via winds and outflows.
@@
In producing these histograms, we have utilized polarimetry data across the range of intensities down to 1000 MJy sr-1. If systematic errors remain after our data cuts, they are unlikely to produce the histogram peaks observed. The scan angles () are sufficiently different from the maxima of the histograms that a systematic scan-synchronous effect is unlikely to produce the results seen here.
To further quantify the locations of the pseudovectors corresponding to the Gaussian peaks of the model fit shown in the left panel of Figure 14 we overlay the vectors that are within 1 of the peaks of the two Gaussian components identified for the full pseudovector histogram distribution (left panel of Figure 14). The pseudovectors corresponding to the -15.6∘ and -78.8∘ model fit components are shown as red and white pseudovectors, respectively, in Figure 15. The locations of the red pseudovectors in the CMZ are observed to largely coincide with molecular clouds (shown in cyan in this figure) and are generally within 0.25∘ in Galactic latitude of the Galactic plane. In particular, the set of red pseudovectors observed at a Galactic Longitude of 359.7∘ traces and coincides with the molecular gas of the “twisted ring” (Molinari et al., 2011). Conversely, the white pseudovectors are generally spread over a larger latitude range.
The tendency of the red pseudovectors to coincide with the molecular clouds supports the idea that a pervasive initially vertical field could have been sheared by the motion of the higher density CMZ clouds. The larger latitude spread of the white pseudovectors supports the idea of a ubiquitous vertical field. The possibility of a ubiquitous field of this nature is further strengthened by the fact that these vertical pseudovectors are not uniquely spatially coincident with the NTFs that are shown in yellow in Figure 15. This finding suggests that the NTFs are illuminated portions of a larger vertical field, rather than being local magnetic field enhancements. A more quantitative analysis of the CMZ-wide magnetic field distribution will be performed in a future work.
| Parameter | All Pseudovectors () | MJy sr-1 |
| 87762 | 45956 | |
| 151778 | 118466 | |
| 172275 | 140364 | |
| -79.6∘1.2∘ | -81.5∘1.2∘ | |
| -15.6∘1.1∘ | -16.0∘1.2∘ | |
| 17.3∘1.2∘ | 16.5∘1.2∘ | |
| 22.8∘1.5∘ | 24.6∘1.6∘ |